OpenStudy (dmcintire7):

Find the minimum sample size n needed to estimate u for the given values of c, o, and E. c=0.98, o=7.4, and E=1 assume the preliminary sample has at least 30 members. N= Round to the nearest whole number @irishboy123

OpenStudy (irishboy123):

what's E?!

OpenStudy (loviy):

Take note of the formula: \[\displaystyle \text{n} = (\frac{\displaystyle \text{z}_{\alpha/2} * \sigma }{ \text{E}} )^2\]

OpenStudy (loviy):

The \[z_{a/2}\] can be found using a z-table.